# From Empirical Observations to Universality: Dynamics of Deep Learning with Inputs Built on Gaussian mixture

### v1

@@ Abstract
This study broadens the scope of theoretical frameworks in deep learning by delving into the dynamics of neural networks with inputs that demonstrate the structural characteristics to Gaussian Mixture (GM). We analyzed how the dynamics of neural networks under GM-structured inputs diverge from the predictions of conventional theories based on simple Gaussian structures. A revelation of our work is the observed convergence of neural network dynamics towards conventional theory even with standardized GM inputs, highlighting an unexpected universality. We found that standardization, especially in conjunction with certain nonlinear functions, plays a critical role in this phenomena. Consequently, despite the complex and varied nature of GM distributions, we demonstrate that neural networks exhibit asymptotic behaviors in line with predictions under simple Gaussian frameworks.
@@ I Introduction
The endeavor to connect practical deep learning applications with theoretical understanding is a growing field of research [1, 2, 3, 4, 5]. Recent studies in this area have begun to study on neural network characteristics under structured nature of input data [6, 7, 8, 9, 10, 11, 12].
@@
The notion of structured data posits that despite the high-dimensional nature of typical datasets (such as MNIST in LeCun [13] has 28x28, CIFAR in Saad and Solla [14] has 3x32x32), these can often be distilled into lower-dimensional representations. This concept is exemplified in the MNIST dataset, where digits, rather than being random pixel assemblies, exhibit structured patterns like lines and circles. This phenomenon of low-dimensional structural features in data has been corroborated by numerous studies [15, 16, 17, 18, 19, 20].
From discussions on the presence of these low-dimensional structural features, considerable research has delved into the intriguing aspects manifested in deep learning when inputs with such characteristics are presented. Particularly, recent studies have found that when inputs inherently exhibit single Gaussian distribution traits, these characteristics are preserved even after passing through the first layer, facilitating a theoretical understanding of deep learning dynamics in simple two-layer models [9, 10, 11].
However, these significant theoretical discoveries fall short of fully mirroring the complexities and heterogeneities of real-world data. Typically, Gaussian mixtures are preferred over single Gaussian distributions for modeling general data characteristics [21, 22, 23], with recent findings suggesting that real-world data often follows Gaussian mixture-like distributions [24, 12]. This gap underscores the necessity for our investigation into Gaussian mixture-based models to capture the complexity of real-world data more accurately. Given this context, our research extends the current theoretical discourse by examining the neural network dynamics when inputs are characterized not by simple Gaussian but by Gaussian mixtures. More specifically, we analyze how neural network dynamics changes as the inherent distribution deviate from simple Gaussian to Gaussian mixture.
Our main findings from investigation into neural network dynamics reflecting Gaussian mixture structural properties are follow:
  • Applying “standardization” to input datasets with inherent Gaussian mixture properties reveals convergence to the predicted dynamics outcomes of existing theories [10].
@@
  • This observed non-divergence is attributed to the distinctive characteristics of nonlinear functions utilized in deep learning network and dataset modeling process, makes deep learning dynamics are predominantly influenced by the distribution’s lower-order cumulants.
The subsequent section, Background II, will present a concise overview of the relevant research. The Methods section III will introduce the Gaussian mixture settings employed in our research and explain the methodology developed to analyze the deviation in dynamics as the distribution changes from a simple Gaussian to a Gaussian mixture. In the Results section IV, we delineate a sequence of experiments showcasing intriguing patterns of convergence, even when the distribution markedly deviates from the typical simple Gaussian distribution. In the Discussion section V, we present a mathematical proof that clarifies the observed phenomena.
Figure 1: Schematic representation of the Hidden Manifold Teacher-Student Model and Discrepancies in Dynamics, Gaussian to Gaussian Mixture Inputs. We investigate the dynamics by calculating the discrepancy measure tLog-RMSE (18, 20), and examining its response to changes in $D_{KL}(P\lVert\mathcal{N})$.
@@ II Background
This section offers a overview of the teacher-student model framework and introduces its evolved variant, the Hidden Manifold Teacher-Student model, as proposed by Goldt et al. [10]. We explain how preceding research efforts have approached the task of understanding theoretical dynamics within the manifold model context.
II.1 Hidden Manifold Teacher-Student Model
The teacher-student model is well-regarded method in the study of high-dimensional problems [25, 26, 4, 5]. This model framework consists of a teacher model, which generates dataset labels, and a student model that learns these labels. Our study specifically zeroes in on the dynamics of a fully-connected two-layer neural network.
The weights of the first and second layers of the teacher model are represented by matrices $\tilde{W}\in\mathbb{R}^{M\times D}$ and $\tilde{v}\in\mathbb{R}^{1\times M}$, respectively. We define the activation function of the teacher model as $\tilde{g}$. Similarly, the weights of the first and second layers of the student model are denoted by $W\in\mathbb{R}^{K\times N}$ and $v\in\mathbb{R}^{1\times K}$, with the activation function represented as $g$.
In the canonical teacher-student model, the input $X\in\mathbb{R}^{N}$ is typically an element-wise i.i.d. from a Gaussian distribution. However, our desired input characteristics are not inherently Gaussian but should reflect intrinsic properties as seen in datasets like the Swiss-roll, which exists on a specific manifold. To embed these intrinsic properties into the input $X$, they utilize a $D$-dimensional vector $C\in\mathbb{R}^{D}$ that follows an element-wise i.i.d. Gaussian distribution. This is achieved through a feature matrix $F\in\mathbb{R}^{D\times N}$ and a nonlinear function $f$ as follows:
$X=f(CF/\sqrt{D})=f(U)\in\mathbb{R}^{N}$ (1)
By modeling the dataset in this manner, the input $X$ intrinsically follows the characteristics of $C$ which is distributed as a Gaussian. Furthermore, the labels generated by the teacher model are derived not directly from $X$, but from $C$, which reflects intrinsic characteristics. In essence, the teacher model generates labels, and the student model learns these labels as follows:
$y=\tilde{g}(C\tilde{W}^{\top}/\sqrt{D})\tilde{v}^{\top},\hat{y}={g}(X{W}^{\top}/\sqrt{N}){v}^{\top}$ (2)
This model, recognizing a hidden structure in lower dimensions, is termed the Hidden Manifold Teacher-Student model [10].
Additionally, for the convenience of subsequent discussions, let’s denote the preactivations of the teacher and student models as $\nu$ and $\lambda$, respectively:
$\nu$ $=C\tilde{W}^{\top}/\sqrt{D}\in\mathbb{R}^{M}$
$\lambda$ $=XW^{\top}/\sqrt{N}=f(U)W^{\top}/\sqrt{N}\in\mathbb{R}^{K}$ (3)
@@ II.1 Hidden Manifold Teacher-Student Model
The notation $\mathsf{M}_{i,j}$ will represent the element located at the $i$-th row and $j$-th column of any given matrix $\mathsf{M}$, and $\mathsf{v}_{i}$ will denote the $i$-th component of any vector $\mathsf{v}$.
Furthermore, in the theoretical analysis to follow, we frequently consider the limit where the dataset dimensions $N$ and intrinsic dimension $D$ become infinitely large ($N\to\infty,D\to\infty$), a scenario often referred to as the thermodynamic limit from the perspective of statistical physics. To maintain consistency with prior theoretical studies, this research also adopts the term thermodynamic limit to describe the $N\to\infty,D\to\infty$ scenario.
II.2 Dynamics of Neural Network are dominant by correlation of function
In this study, we update weights through a simple scaled stochastic gradient descent (SGD) with a batch size of one, under quadratic loss $\mathcal{L}=1/2(\hat{y}-y)^{2}$:
$W_{k,i}$ $:=W_{k,i}-\frac{\eta}{\sqrt{N}}v_{k}(\hat{y}-y)g^{\prime}\left(\lambda_{k}\right)f(U_{i})$
$v_{k}$ $:=v_{k}-\frac{\eta}{N}g\left(\lambda_{k}\right)(\hat{y}-y)$
To illustrate a straightforward example, let’s examine the dynamics through the explicit ordinary differential equation (ODE) form of the second weight, $v$. Defining the normalized number of steps as $t=1/N$ in the thermodynamic limit $N\to\infty$, which can be interpreted as a continuous time-like variable. Consequently, $v_{k}$ satisfies the following ODE.
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\tilde{v}_{n}I_{2}(k,m)-\sum^{K}_{j}v_{j}I_{2}(k,j)\right]$ (4)
where $I_{2}(k,m)=\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]$ and $I_{2}(k,j)=\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]$ represent the correlations of function.
Using a similar approach for $v$, we can derive the dynamics of our teacher-student model in the form of ODEs, as detailed in the Appendix B. The dynamics are predominantly influenced by the correlations of specific functions, like $\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})],\mathbb{E}[g(\lambda_{k})g(\lambda_{j})]$. To calculate these function correlation values, we requires information on the underlying distribution of $\{\lambda,\nu\}$.
@@ II.3 The Gaussian Equivalence Property
Previous analyses have delved into understanding the distribution of preactivations $\{\lambda,\nu\}$ under certain assumptions. To summarize the findings, $\{\lambda,\nu\}$ adhere to a Gaussian distribution characterized by a specific covariance matrix.
To provide a simplified derivation, we first explore how the function correlation approximately follows. Suppose random variables from a joint Gaussian distribution $\{x_{1},x_{2}\}$ are weakly correlated ($\mathbb{E}[x_{1}x_{2}]\sim\mathcal{O}(\epsilon)$), and arbitrary functions $u,v$ are regular enough to guarantee the existence of an expectation value, then the following lemma holds:
@@
$\mathbb{E}[u(x_{1})v(x_{2})]\approx\mathbb{E}[u(x_{1})]\mathbb{E}[v(x_{2})]+\mathcal{O}(\epsilon)$ (5)
If the feature matrix $F$ and the student weight matrix $W$ are sufficiently bounded, fulfilling the following assumption:
@@
For all $p,q\geq 1$ and any indices $k_{1},\cdots,k_{p},r_{1},\cdots,r_{p}$:
$\cfrac{1}{\sqrt{N}}\sum_{i}W_{k_{1},i}\cdots W_{k_{p},i}\times F_{r_{1},i}\cdots F_{r_{q},i}=\mathcal{O}(1)$ (6)
we can approximately calculate the covariances of $\{\lambda,\nu\}$, i.e., $\mathbb{E}[\lambda\lambda],\mathbb{E}[\nu\lambda],\mathbb{E}[\nu\nu]$. By referring to the definition of $\{\lambda,\nu\}$ and employing the lemma II.1 to decompose each function correlation, and then sorting out the terms that vanish in the thermodynamic limit ($N\to\infty,D\to\infty$) according to the bounded assumption (Appendix A.3).
This analysis enables the derivation of the asymptotic form of all covariance matrices, where higher-order correlation vanish in the thermodynamic limit [10]. Consequently, the preactivations follow a Gaussian distribution, a result summarized as the Gaussian Equivalence Property (GEP).
@@
In the thermodynamic limit ($N\to\infty$, $D\to\infty$), with finite $K$, $M$, $D/N$, and under the assumption II.2, if the $C$ follows a normal Gaussian distribution $\mathcal{N}(0,I)$, then $\{\lambda,\nu\}$ conform to $K+M$ jointly Gaussian variables. This means that statistics involving $\{\lambda,\nu\}$ are entirely represented by their mean and covariance.
Just as the dynamics of $v$ significantly depend on the distribution characteristics of $\{\lambda,\nu\}$, the neural network’s dynamics can be analyzed through these distribution characteristics (Appendix B). This property II.3 allows for an understanding of the student and teacher models’ dynamics, via the mean and covariance of the joint Gaussian distribution of preactivations. A comprehensive derivation and discussion are provided in the appendix A and B.
For convenience, let’s redefine $\bar{\lambda}_{k}$ as:
$\bar{\lambda}_{k}=\cfrac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}(f(U_{i})-\mathbb{E}_{u\sim\mathcal{N}(0,1)}[f(u)])$ (7)
$\bar{\lambda}_{k}$ also follows a jointly Gaussian distribution, and its expectation value satisfies $\mathbb{E}[\bar{\lambda}_{k}]=0$.
Consequently, the new distribution $\{\bar{\lambda},\nu\}$ follows a more straightforward distribution with the mean
$\mathbb{E}\left[\bar{\lambda}_{k}\right]=\mathbb{E}\left[\nu_{m}\right]=0$ (8)
and the covariance
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\bar{\lambda}}_{k}{\bar{\lambda}}_{\ell}\right]=\left(c-a^{2}-b^{2}\right)\Omega_{k,\ell}+b^{2}\Sigma_{k,\ell}$ (9)
$R_{k,m}$ $\equiv\mathbb{E}\left[\bar{\lambda}_{k}\nu_{m}\right]=b\frac{1}{D}\sum_{r=1}^{D}S_{k,r}\tilde{W}_{m,r}$ (10)
$T_{m,n}$ $\equiv\mathbb{E}\left[\nu_{m}\nu_{n}\right]=\frac{1}{D}\sum_{r=1}^{D}\tilde{W}_{m,r}\tilde{W}_{n,r}.$ (11)
Here, $a$, $b$, and $c$ represent the statistical properties of the nonlinear function $f$, used in the transformation of student model inputs $X$, $X=f(CF/\sqrt{D})$,
$a=\mathbb{E}[f(u)],\quad b=\mathbb{E}[uf(u)],\quad c=\mathbb{E}[f(u)^{2}]$ (12)
@@
$S_{k,r}$ $\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}F_{r,i}$ (13)
$\Omega_{k,\ell}$ $\equiv\frac{1}{N}\sum_{i=1}^{N}W_{k,i}W_{\ell,i}$ (14)
$\Sigma_{k,\ell}$ $\equiv\frac{1}{D}\sum_{r=1}^{D}S_{k,r}S_{\ell,r}$ (15)
These defined covariances capture essential characteristics inherent to the teacher-student model dynamics. Given that the weights of the student model, $v$ and $W$, evolve under the stochastic gradient descent (SGD) dynamics, these values, and consequently the covariances $Q$, $R$, $T$, inherently depend on the progression of training steps (time).
The student model learns by attempting to emulate the teacher model’s outputs. Each covariance matrix holds a distinct significance in relation to the dynamics of the model. Specifically, $R$ signifies the correlation between the preactivations of the student and teacher models, reflecting the student model’s accuracy in mirroring the teacher. Conversely, $Q$ relates to the correlation among the student model’s own preactivations, implying the dynamics of the student model’s first layer. While $T$ remains constant throughout the learning process and thus stands apart from $Q$ and $R$, it serves as a mirror to the teacher model’s inherent characteristics.
To summarize, within the context of a hidden manifold teacher-student model characterized by simple Gaussian distribution properties, the learning dynamics of the student model are primarily influenced by terms like $\mathbb{E}[g(\lambda)\tilde{g}(\nu)]$. Such terms, which depend on the distributional properties of $\{\lambda,\nu\}$, can be calculated once distribution determined. Under certain assumptions, the GEP elucidates that the preactivations follow a Gaussian distribution. Consequently, this allows us to analytically dissect the dynamics of the student model.

### v2

# Gaussian Universality in Neural Network Dynamics with Generalized Structured Input Distributions

@@ Abstract
Bridging the gap between the practical performance of deep learning and its theoretical foundations often involves analyzing neural networks through stochastic gradient descent (SGD). Expanding on previous research that focused on modeling structured inputs under a simple Gaussian setting, we analyze the behavior of a deep learning system trained on inputs modeled as Gaussian mixtures to better simulate more general structured inputs. Through empirical analysis and theoretical investigation, we demonstrate that under certain standardization schemes, the deep learning model converges toward Gaussian setting behavior, even when the input data follow more complex or real-world distributions. This finding exhibits a form of universality in which diverse structured distributions yield results consistent with Gaussian assumptions, which can support the theoretical understanding of deep learning models.
@@ I Introduction
The study of artificial neural networks through the lens of statistical physics has a well-established history. Neural networks trained on samples from a distribution have traditionally been analyzed as optimization problems within complex systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These problems are generally classified based on the learning process, whether the entire dataset or a subset is used per iteration, as in gradient descent and batch gradient descent, respectively, or whether a single sample is used per iteration, as in stochastic gradient descent (SGD, also known as on-line learning).
When the entire dataset is used, one can fix the dataset size and interpret the neural network and corresponding loss function as analogous to a spin system and potential energy. The replica method can then be employed to investigate the system. This approach has provided successful interpretations of simple one- or two-layer perceptrons during the early development of neural networks [2, 3, 4, 5], and more recently has demonstrated applicability to more advanced settings including generative models [12, 13, 14].
On the other hand, when considering SGD, where a single dataset is used per iteration, the neural network variables are updated to optimize the loss for each sample. The behavior of neural networks under such independent random sampling can be described without the need for the replica method. By analyzing the equations of motion derived from this approach, it is possible to track generalization error and the time evolution of specific weights in the neural network [6, 7, 8, 9, 10, 11].
@@
Recently, the composition of the input dataset has become a crucial consideration. Studies have shifted focus toward understanding neural network behavior under structured input data [15, 16, 17, 18, 19, 20, 21]. The notion of structured data posits that despite the high-dimensional nature of typical datasets (such as MNIST in Deng [22] with 28$\times$28 dimensions and CIFAR in Krizhevsky [23] with 3$\times$32$\times$32 dimensions), they can often be distilled into lower-dimensional representations. For instance, in the MNIST dataset, the digits display structured patterns like lines and curves rather than random pixel arrangements. This characteristic of low-dimensional structural features in data has been discussed in numerous studies [24, 25, 26, 27, 28, 29].
Motivated by the presence of these low-dimensional structures, recent research has explored the impact of processing such structured inputs on deep learning. Specifically, theoretical analyses have modeled inputs with a simple Gaussian distribution [18, 19, 20]; however, such theoretical approaches are still limited in their ability to capture the complexities of real-world data, which are often better represented by Gaussian mixtures rather than simple Gaussian distributions [30, 31, 32].
Our study aims to extend the simple Gaussian framework by examining neural network dynamics when inputs are characterized by Gaussian mixtures in low dimensions. We analyze how neural network behavior changes as the underlying distribution shifts from simple Gaussian to a Gaussian mixture.
Figure 1: Illustration of the hidden manifold model (Goldt et al. [19]) and our experimental scheme using Gaussian and Gaussian mixture inputs. The term ODE dynamics refers to the outcomes from ordinary differential equation (ODE) simulations, which align with SGD under a simple Gaussian input. In contrast, SGD dynamics describes the results obtained from running SGD with $C\sim\mathcal{P}$.
Our key findings are as follows:
  • Standardization of input datasets modeled by Gaussian mixtures (or general distributions) in low dimensions results in convergence with the dynamics observed under Gaussian inputs.
@@
  • The observed convergence is largely due to the nonlinear functions in deep learning models, which make the network dynamics predominantly sensitive to the distribution’s lower-order cumulants.
This study is organized as follows. In Section II we provide a brief background on relevant studies, and in Section III we describe the Gaussian mixture settings and experimental conditions. In Section IV, we present the experimental results showcasing the patterns of convergence. Then in Section V we offer mathematical proof of the observed phenomena and applications with pseudo-real datasets. Finally, we conclude our study in Section VI.
@@ II Background
This section offers an overview of the teacher-student model framework and one of its evolved variants, the hidden manifold model proposed by Goldt et al. [19].
II.1 Hidden Manifold Model
The teacher-student model is a well-regarded method in the study of high-dimensional problems [33, 34, 35, 36]. This model framework consists of a teacher model that generates dataset labels and a student model that learns the labels. In a two-layer neural network, the weights of the first and second layers of the teacher model are represented by matrices $\widetilde{W}\in\mathbb{R}^{M\times D}$ and $\widetilde{v}\in\mathbb{R}^{1\times M}$, respectively. We define the activation function of the teacher model as $\widetilde{g}$. Similarly, the weights of the first and second layers of the student model are denoted by $W\in\mathbb{R}^{K\times N}$ and $v\in\mathbb{R}^{1\times K}$, with the activation function represented as $g$.
In the canonical teacher-student model, the input $X\in\mathbb{R}^{N}$ is typically element-wise i.i.d. from a Gaussian distribution. But here, we aim for input characteristics that reflect intrinsic structural properties rather than being merely extrinsically Gaussian.
To embed these intrinsic properties into the input $X$, Goldt et al. [19] utilized a $D$-dimensional vector, $C\in\mathbb{R}^{D}$, that follows an element-wise i.i.d. Gaussian distribution. This is achieved through the feature matrix $F\in\mathbb{R}^{D\times N}$ and nonlinear function $f$ as follows:
$X=f(CF/\sqrt{D})=f(U)\in\mathbb{R}^{N}.$ (1)
By modeling the dataset in this manner, the input $X$ intrinsically reflects the characteristics of $C$, which is distributed as Gaussian. Furthermore, the labels generated by the teacher model are derived not directly from $X$ but from $C$, which reflects the dominance of the intrinsic characteristics in the true label.
In summary, the teacher-student model results can be expressed as follows:
$y=\widetilde{g}(C\widetilde{W}^{\top}/\sqrt{D})\widetilde{v}^{\top},\hat{y}={g}(X{W}^{\top}/\sqrt{N}){v}^{\top}.$ (2)
This model, recognizing a hidden structure in lower dimensions, is termed the hidden manifold model [19]; Fig. 1 shows a schematic of the model along with our experimental scheme (see Section IV). For the convenience of subsequent discussions, we denote the preactivations of the teacher and student models as $\nu$ and $\lambda$, respectively:
$\nu$ $=C\widetilde{W}^{\top}/\sqrt{D}\in\mathbb{R}^{M},$
$\lambda$ $=XW^{\top}/\sqrt{N}=f(U)W^{\top}/\sqrt{N}\in\mathbb{R}^{K}.$ (3)
@@ II.1 Hidden Manifold Model
The notation $\mathsf{M}_{i,j}$ represents the element located in the $i$-th row and $j$-th column of any given matrix $\mathsf{M}$, and $\mathsf{v}_{i}$ denotes the $i$-th component of any vector $\mathsf{v}$.
Below, we frequently consider an infinitely large dataset dimension $N$ and intrinsic dimension $D$ ($N\to\infty,D\to\infty$), a scenario often referred to as the thermodynamic limit from the perspective of statistical physics. To maintain consistency with prior studies, this research also adopts the term thermodynamic limit to describe the $N\to\infty$, $D\to\infty$ scenario.
II.2 Dynamics of Neural Network Weights
In this study, we use the Goldt et al. [19] approach to update the student model weights through a scaled SGD under quadratic loss $\mathcal{L}=1/2(\hat{y}-y)^{2}$,
$W_{k,i}$ $:=W_{k,i}-\frac{\eta}{\sqrt{N}}v_{k}(\hat{y}-y)g^{\prime}\left(\lambda_{k}\right)f(U_{i}),$ (4)
$v_{k}$ $:=v_{k}-\frac{\eta}{N}g\left(\lambda_{k}\right)(\hat{y}-y).$
By defining the normalized number of steps as $t=1/N$ in the thermodynamic limit $N\to\infty$, which can be interpreted as a continuous time-like variable, the updates transform into ordinary differential equations (ODEs).
For example, $v_{k}$ satisfies the following ODE:
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\widetilde{v}_{n}I_{2}(k,m)-\sum^{K}_{j}v_{j}I_{2}(k,j)\right],$ (5)
where $I_{2}(k,m)=\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{m})]$ and $I_{2}(k,j)=\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]$ represent the correlations of functions.
Using a similar approach for $v$, we can derive the dynamics of our teacher-student model in the form of ODEs, as detailed in Appendix A. The dynamics are predominantly influenced by the correlations of specific functions, such as $I_{2}(k,m)$, and $I_{2}(k,j)$. To calculate these correlations of functions, referred to here as the function correlation, we require information on the underlying distribution of $\{\lambda,\nu\}$.
@@ II.3 The Gaussian Equivalence Property
Earlier works have investigated the distribution of preactivations $\{\lambda,\nu\}$ under certain assumptions. To summarize the previous findings, $\{\lambda,\nu\}$ follow a Gaussian distribution characterized by a specific covariance matrix [19].
To provide a simplified derivation, we first explore how the function correlation is approximated, where we can decompose it under weakly correlated conditions.
@@
If random variables from a joint Gaussian distribution $\{x_{1},x_{2}\}$ are weakly correlated, i.e., $\mathbb{E}[x_{1}x_{2}]\sim\mathcal{O}(\epsilon)$, and arbitrary functions $u,v$ are sufficiently regular to guarantee the existence of expectation values, then
$\mathbb{E}[u(x_{1})v(x_{2})]\approx\mathbb{E}[u(x_{1})]\mathbb{E}[v(x_{2})]+\mathcal{O}(\epsilon).$ (6)
Additionally, the weights $W,\widetilde{W}$ and feature matrix $F$ do not significantly alter the input properties, having elements in $\mathcal{O}(1)$ scale. This allows us to make the following bounded assumption.
@@
For all $p,q\geq 1$ and any indices $k_{1},\cdots,k_{p},r_{1},\cdots,r_{q}$:
$\cfrac{1}{\sqrt{N}}\sum_{i}W_{k_{1},i}\cdots W_{k_{p},i}\times F_{r_{1},i}\cdots F_{r_{q},i}=\mathcal{O}(1),$ (7)
with the $q$ and $p$ distinct.
Under these conditions, we can approximately calculate the covariances of $\{\lambda,\nu\}$, i.e., $\mathbb{E}\left[\lambda\lambda\right],\mathbb{E}[\nu\lambda],\mathbb{E}[\nu\nu]$. This analysis enables the derivation of the asymptotic form of all covariance matrices of preactivations $\{\lambda,\nu\}$, where higher-order correlations vanish in the thermodynamic limit [19].
Consequently, the preactivations follow a Gaussian distribution, a result summarized as the Gaussian equivalence property (GEP).
@@
In the thermodynamic limit ($N\to\infty$, $D\to\infty$), with finite $K$, $M$, $D/N$, and under Assumption II.2, if $C$ follows a standard Gaussian distribution $\mathcal{N}(0,I)$, then $\{\lambda,\nu\}$ are jointly Gaussian variables of dimension $K+M$. This means that the statistics involving $\{\lambda,\nu\}$ are entirely represented by their mean and covariance.
Property II.3 states that the preactivation distribution $\{\lambda,\nu\}$ follows a joint Gaussian distribution, if the student weights satisfy Assumption II.2. This property allows us to understand the student model’s dynamics (e.g., student weight, generalization error) via the mean and covariance of the joint Gaussian distribution of preactivations. For example, the student weight dynamics (Eq. (5)) are tractable by calculating $I_{2}$ under the $\{\lambda,\nu\}$ distribution. Detailed derivations are provided in Appendix A.
Before expressing the covariance matrix, for convenience we redefine $\overline{\lambda}_{k}$ as
$\overline{\lambda}_{k}=\cfrac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}(f(U_{i})-\mathbb{E}_{u\sim\mathcal{N}(0,1)}[f(u)]).$ (8)
$\overline{\lambda}_{k}$ also follows a jointly Gaussian distribution with $\mathbb{E}[\overline{\lambda}_{k}]=0$. Consequently, the new distribution $\{\overline{\lambda},\nu\}$ has means of
$\mathbb{E}\left[\overline{\lambda}_{k}\right]=\mathbb{E}\left[\nu_{m}\right]=0,$ (9)
and covariances as follows:
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\overline{\lambda}}_{k}{\overline{\lambda}}_{\ell}\right]=\left(c-a^{2}-b^{2}\right)\Omega_{k,\ell}+b^{2}\Sigma_{k,\ell},$ (10)
$R_{k,m}$ $\equiv\mathbb{E}\left[\overline{\lambda}_{k}\nu_{m}\right]=b\frac{1}{D}\sum_{r=1}^{D}S_{k,r}\widetilde{W}_{m,r},$ (11)
$T_{m,n}$ $\equiv\mathbb{E}\left[\nu_{m}\nu_{n}\right]=\frac{1}{D}\sum_{r=1}^{D}\widetilde{W}_{m,r}\widetilde{W}_{n,r}.$ (12)
Here, $a$, $b$, and $c$ represent the statistical properties of the nonlinear function $f$, used in the transformation of student model inputs $X$, $X=f(CF/\sqrt{D})$, given as
$a=\mathbb{E}[f(u)],\quad b=\mathbb{E}[uf(u)],\quad c=\mathbb{E}[f(u)^{2}],$ (13)
@@
$S_{k,r}$ $\equiv\frac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}F_{r,i},$ (14)
$\Omega_{k,\ell}$ $\equiv\frac{1}{N}\sum_{i=1}^{N}W_{k,i}W_{\ell,i},$ (15)
$\Sigma_{k,\ell}$ $\equiv\frac{1}{D}\sum_{r=1}^{D}S_{k,r}S_{\ell,r}.$ (16)
For compact notation, we focus on the symmetric nonlinear function $f$ that satisfies $a=\mathbb{E}[f(u)]=0$. These defined covariances capture the essential characteristics of the teacher-student model dynamics.
The student model learns by attempting to emulate the teacher model outputs. Each covariance matrix holds a distinct meaning in relation to the dynamics of the model. The matrix $Q$ relates to the correlation among the student model’s own preactivations, implying the dynamics of the first layer of the student model. The matrix $R$ relates to the correlation between the preactivations of the student and teacher models, reflecting the student model’s accuracy in mirroring the teacher. The matrix $T$ remains constant throughout the learning process, serving as a mirror of the teacher model’s inherent characteristics.
To summarize, within the context of the hidden manifold model characterized by a simple Gaussian distribution, the learning dynamics of the student model are primarily influenced by the function correlation terms. Such terms, which depend on the distributional properties of $\{\lambda,\nu\}$, can be calculated once the distribution is determined. Under certain assumptions, the GEP shows that the preactivations $\{\lambda,\nu\}$ follow a Gaussian distribution, allowing us to analytically dissect the dynamics of the student model.